In coming to a stop, a car leaves skid marks 65 mPLEASE HELP!!
Answer #1

distance covered from skid mark, S = 65m

deceleration (means negative acceleration) , a = - 4 m/s2

Vf (final speed) of car = 0 m/s

Vi ( initial speed of car before braking) = ?

from equation of motion:

Vf2 - Vi2 = 2*a*S

0 - Vi2 = 2*( -4)* 65

Vi = 22.80 m/s